OpenStudy (anonymous):

If x varies inversely as y and x = 48 when y = 4, find y when x = 3. A. y = 16 B. y = 12 C. y = −4 D. y = 64

OpenStudy (aroub):

Because it varies inversely so it's, xy=k, we just substitute the values that are given in the equation to find k: xy=k 48(4)=k 192=k xy=k 3(y)=192 y=64

OpenStudy (aroub):

And you can also use this way: x1*y1=x2*y2 (i think it's easier)

OpenStudy (anonymous):

how about this one?

OpenStudy (anonymous):

If x varies inversely as y and x = 10 when y = 20, find y when x = 25. A. y = 8 B. y = 5 C. y = −10 D. y = 4

OpenStudy (aroub):

What about you try? :)

OpenStudy (anonymous):

i've already did and couldnt get it and i've asked someone on here and they couldnt figure it out

OpenStudy (aroub):

O.o Okay I'll do it but using the second way. x1*y1=x2*y2 10(20)=25(y) 200=25y y=8